# TCIP: Threshold-Controlled Iterative Pyramid Network for Deformable Medical Image Registration

### v1

@@ 4.2 Generalizability of FERM and TCI
Dice (%) MSE Time (s) Dice (%) MSE Time (s)
IIRP 65.9 4.16 0.12 72.2 0.62 0.10
IIRP + TCI 66.2 3.56 0.14 73.0 0.61 0.14
@@ 4.3 Ablation Studies
We further conduct ablation studies to assess the effectiveness of the two key FERM modules, namely FFB and SEB. The performance of variants w/o FFB and/or SEB are shown in Table 5. As shown, the variant with neither FFB and SEB yields the lowest performance, indicating that relying solely on DeF is insufficient to capture complex deformation patterns, leading to higher registration errors. Furthermore, removing either FFB or SEB degrades performance relative to the intact FERM, demonstrating that the integration of them both enhances the network’s ability to accurately capture fine structural details and suppress irrelevant features, thereby significantly improving registration accuracy.
Table 5: Ablation study on different FERM modules
FFB SEB Mindboggle LPBA IXI
Dice(%) MSE Dice(s) MSE Dice(s) MSE
@@
65.9 4.02 72.7 0.71 80.4 0.24
66.4 3.97 73.3 0.61 80.5 0.23
In addition, we investigate the performance of four TCI configurations on the IXI dataset. Specifically, TCI-1 performs only convergence evaluation, TCI-2 performs only stability evaluation, TCI-3 evaluates convergence before stability, and TCI-4 evaluates stability before convergence (the configuration adopted in TCIP). Table 6 shows that TCI-4 performs best, showcasing the appropriateness of TCI design.
@@ 5 Conclusion
We propose the deformable image registration network TCIP, which employs FERM to extract anatomical structure details and learn to suppress irrelevant features. Additionally, we introduce the dual-stage TCI strategy to determine the number of iterations. The results on four datasets demonstrate the effectiveness of TCIP. Generalizability and ablation studies validate the effectiveness of FERM and TCI. Currently, we extend the original 2D SEB into a 3D variant to suppress irrelevant features. Going forward, we believe further improvement is achievable by extending SEB to incorporate local importance pooling to emphasize key regions rather than relying solely on global average pooling.
@@ Appendix A Datasets and Implementation Details
Datasets: To comprehensively assess the performance of the proposed TCIP, we compare it against different registration models on three public MRI datasets, namely LPBA (Shattuck et al. 2008), Mindboggle (Klein and Tourville 2012), IXI (Avants et al. 2009), and one abdomen CT dataset FLARE (Ma et al. 2022). We preprocess these four datasets as follows. The LPBA dataset contains 40 T1-weighted MRI volumes, each annotated with 54 manually labeled Regions of Interest (ROIs). All volumes have been rigorously pre-aligned to the MNI 305. Specifically, each volume is cropped to $160\times 192\times 160$ voxels with an isotropic voxel size of 1 mm. We use 31 volumes for training and 9 for testing, following the prior study (Wang, Ni, and Wang 2024). For the Mindboggle dataset, we choose the NKI-RS-22 and NKI-TRT-20 sub-datasets from Mindboggle101 as the training set and OASIS-TRT-20 as the test set, following the prior study (Ma et al. 2024). We crop each volume to $160\times 192\times 160$ with an isotropic voxel size of 1 mm, while all images are pre-aligned with the MNI 152 template space. The IXI dataset contains the original brain MRI images provided by (Fischl 2012) and their corresponding 3D volume data after affine registration. All volumes are resampled to a voxel size of $80\times 96\times 112$ (isotropic voxel size of 2 mm), and for training, we use 407 3D volumes (total $407\times 406$ paired samples), and for testing, 58 volumes (total $58\times 57$ paired samples), following the prior study (Wang, Ni, and Wang 2024). The FLARE dataset consists of 361 scans annotated with four organ labels, namely liver, kidney, spleen, and pancreas. Each volume is cropped to $160\times 192\times 160$ voxels with an isotropic voxel size of 2.5 mm. We allocate 341 volumes for training and 20 for testing following the (Ma et al. 2024).
@@
Metrics: To comprehensively assess the performance of various registration models, we adopt multiple quantitative metrics. The Dice score (Dice 1945) measures region‑wise overlap, while the 95% maximum Hausdorff Distance (HD95) (Huttenlocher, Klanderman, and Rucklidge 1993) and the Average Symmetric Surface Distance (ASSD) (Taha and Hanbury 2015) quantify the similarity of the regional profiles. The quality of the estimated deformation field $\phi$ is assessed by the percentage of voxels with non-positive Jacobian determinant $|{J}_{\phi}|_{<0}$↓ (i.e., folded voxels). In addition, we report Mean Squared Error (MSE), Inference Time (s), and model size (the number of parameters) for a comprehensive performance analysis.
@@
Implementation Details: Our Model is implemented using Pytorch and trained on an NVIDIA GeForce RTX 4090 GPU. Following the prior study (Ma et al. 2024), we normalize all images to the range $[0,1]$, optimize using the ADAM optimizer (Kingma 2014) at a learning rate of $1\times 10^{-4}$, and set the size of volumetric patch $n$, hyperparameter $\lambda$, and batch size to 9, 1, and 1, respectively. In FERM, we adopt a reduction ratio $r=16$. In TCI, we set the window size $t$ and the maximum number of iterations $k_{max}$ to 3 and 10, respectively.
Appendix B Performance Comparison on HD95, ASSD, and MSE Metrics
@@
Table B: Performance comparison of different registration models on the Mindboggle, LPBA, and IXI datasets evaluated by HD95, ASSD, and MSE metrics
@@
HD95↓ ASSD↓ MSE ($10^{-3}$) HD95↓ ASSD↓ MSE ($10^{-3}$) HD95↓ ASSD↓ MSE ($10^{-3}$)
U-Net
VM (CVPR’18) 5.683 1.726 7.449 6.557 2.089 2.293 1.460 0.584 0.727
@@
TCIP (ours) 4.909 1.375 3.969 5.577 1.657 0.614 1.229 0.470 0.230
In Table B, we present the performance of different registration models on the Mindboggle, LPBA, and IXI datasets, evaluated by the HD95, ASSD, and MSE metrics. The experimental results demonstrate that TCIP outperforms the other registration models in most scenarios and achieves SOTA results on at least two metrics for each dataset.
Appendix C Quantitative Analysis on Different Anatomical Structures
Figure C.1: Performance comparison of different models in seven regions of the LPBA dataset.
Figure C.2: Performance comparison of different models in seven regions of the Mindboggle dataset.
Figure C.3: Performance comparison of different models in thirty regions of the IXI dataset.
Following the prior study (Wang, Ni, and Wang 2024), we evaluate TCIP’s registration performance on selected regions in both hemispheres to assess its accuracy across diverse anatomical structures. Specifically, Figure C.1 presents the Dice scores for seven brain regions in the LPBA dataset, including the frontal, parietal, temporal, occipital, cingulate, hippocampus, and putamen. Figure C.2 presents the Dice scores for 7 brain regions in the Mindboggle dataset, including the frontal, parietal, temporal, occipital, third ventricle, fourth ventricle, and brainstem. Figure C.3 presents the Dice score across 30 regions of the IXI dataset. All these experimental results demonstrate that TCIP outperforms existing registration models.
Appendix D Effectiveness of TCI at Each Decoding Layer
Table D: Ablation study on the effectiveness of TCI on various decoding layers
@@ Appendix D Effectiveness of TCI at Each Decoding Layer
65.9 73.0 80.1
66.1 73.2 80.2
66.3 73.3 80.4
66.4 73.3 80.5
In this section, we conduct ablation studies to assess the effectiveness of TCI at each decoding layers. Omitting TCI at a given layer fixes the number of iterations to one, thereby preventing iterative refinement of the deformation field. As shown in Table D, progressively enabling TCI across additional layers steadily improves TCIP’s performance. Notably, applying TCI at higher-resolution layers (e.g., Layer 1) yields larger performance gains than at lower-resolution layers (e.g., Layer 4). A plausible reason for this is that higher-resolution feature maps more closely match the original image dimensions, preserving fine-grained structural details. Consequently, determining the number of iterations for higher-resolution layers yields greater performance improvements.
Appendix E Impact of Varying Values of Thresholds
Table E: Ablation study on thresholds $\delta_{s}$ and $\delta_{c}$ values
@@ Appendix E Impact of Varying Values of Thresholds
$0.01$ 65.9 66.1 73.1 73.2 80.4 80.4
$0.005$ 66.2 66.4 73.2 73.3 80.5 80.5
$0.001$ 66.2 66.2 73.1 73.1 80.4 80.2
Figure E: Statistics of the number of iterations across different decoding layers.
In this section, we investigate the sensitivity of the predetermined thresholds $\delta_{s}$ and $\delta_{c}$. As shown in Table E, these two thresholds exhibit low sensitivity to variation and have minimal impact on the overall performance of TCIP. Nonetheless, we choose the middle range values, i.e., $\delta_{s}=0.005$ and $\delta_{c}=0.005$, which yield the best performance
In addition, in Figure E, we present the number of iterations per decoding layer determined by TCI on the IXI dataset. The experimental results demonstrate that three or four iterations suffice for convergence in most cases. Hence, our predefined maximum number of iterations $K_{max}=10$ is reasonable.
Appendix F Impact of Different TCI Similarity Metrics
In this section, we examine the effect of three similarity metrics, namely Normalized Cross-Correlation (NCC), Mean Absolute Error (MAE), and Mean Squared Error (MSE), on the performance of TCI. As shown in Table F, NCC delivers the best results. Hence, we adopt NCC as the similarity metric for all experiments in this paper.
Table F: Ablation study on the effectiveness of different similarity metrics adopted in TCI
@@ Appendix F Impact of Different TCI Similarity Metrics
Appendix G Impact of Different Sliding Window Sizes
Table G: Ablation study on the different window sizes
IXI
$t$ Dice HD95 ASSD MSE time
3 80.4 1.23 0.48 0.24 0.09
4 80.9 1.22 0.46 0.21 0.15
5 80.7 1.27 0.49 0.23 0.29
Mindboggle
$t$ Dice HD95 ASSD MSE time
3 66.4 4.91 1.38 3.97 0.16
4 66.1 4.93 1.38 3.89 0.22
5 65.8 4.94 1.39 4.06 0.35
In Table G, we present the effect of window size $t$ on the performance of TCIP. On the IXI dataset, increasing $t$ from 3 to 4 yields marginal performance gains but incurs longer inference time. Increasing $t$ further to 5 degrades performance across all metrics, suggesting that $t=4$ provides sufficient iterative refinement without incurring excessive iterations. On the Mindboggle dataset, setting $t=3$ outperforms $t=4$ on most metrics, likely because $t=4$ induces excessive iterations in TCI’s deformation refinement. Therefore, we set $t=3$ for all experiments in this paper.

### v2

@@ 4.2 Generalizability of FARM and TCI
Dice(%) MSE↓ Time(s)↓ Dice(%) MSE↓ Time(s)↓
IIRP 65.8 4.16 0.12 72.2 0.62 0.10
[…]
To demonstrate the generalizability of FARM and TCI, we integrate each into existing registration models and report the results in Tables 3 and 4, respectively. Table 3 shows that replacing the decoder with FARM yields consistent performance gains across all three pyramid models. This is because FARM adaptively emphasizes anatomically relevant features while suppressing less informative ones, thereby mitigating error accumulation during coarse-to-fine decoding. Furthermore, Table 4 demonstrates that substituting IIRP’s original single-stage iterative strategy with TCI yields performance improvement, underscoring the advantage of TCI in determining the number of optimization iterations dynamically. To provide in-depth details of this gain, we present the number of iterations determined by TCI and by the single-stage strategy across the four decoding levels of IIRP in Appendix F. Overall, these results indicate that TCI can alleviate premature termination.
@@ 4.3 Ablation Studies
We further conduct ablation studies to assess the effectiveness of all components in FARM. Table 5 reports ablations by enabling/disabling SEB, SAB, and RP in different combinations. Because spatial attention relies on channel-wise aggregation and should be applied on well-calibrated channel representations, we do not evaluate the variant swapping the order of SEB and SAB. It is not surprising that the model performs worst when SEB, SAB, and RP are all removed (relying solely on a single convolutional layer ${Cl}_{3}$). Moreover, removing any individual module or pathway also degrades performance comparing to the intact FARM. These results suggest that SEB, SAB, and RP act synergistically, enhancing the capture of subtle anatomical structures while suppressing irrelevant features, thereby improving registration accuracy both individually and holistically.
Table 5: Ablation study on different FARM modules
SEB SAB RP Mindboggle LPBA IXI
Dice(%)↑ MSE↓ Dice(%)↑ MSE↓ Dice(%)↑ MSE↓
@@
66.2 4.02 73.4 0.71 80.4 0.24
66.1 4.00 73.3 0.68 80.3 0.24
67.5 3.98 73.5 0.65 80.4 0.23
67.9 3.97 73.7 0.61 80.6 0.23
In addition, we investigate the performance of four TCI configurations on the IXI dataset. Specifically, TCI-1 performs only convergence evaluation, TCI-2 performs only stability evaluation, TCI-3 evaluates convergence before stability, and TCI-4 evaluates stability before convergence (the configuration adopted in iPEAR). Table 6 shows that TCI-4 performs best, showcasing the appropriateness of the proposed TCI design.
@@ 5 Conclusion
We propose the deformable image registration network iPEAR, which employs FARM to extract anatomical structure details and learn to suppress irrelevant features. Additionally, we introduce the dual-stage TCI strategy to determine the number of optimization iterations. The results on four datasets demonstrate the effectiveness of iPEAR. Generalizability and ablation studies validate the effectiveness of FARM and TCI. Currently, we combine channel attention and spatial attention to suppress irrelevant features. Going forward, we believe further improvement is achievable by extending this to a multi-head attention mechanism to enhance the model’s capability in feature representation.
@@ Appendix A Datasets and Implementation Details
Datasets: Consistent with prior studies (Wang and Wang, 2025; Wang et al., 2024), iPEAR focuses on single-modality registration. Accordingly, to comprehensively assess the performance of the proposed iPEAR, we compare it against different registration models on three public MRI datasets, namely Mindboggle (Klein and Tourville, 2012), LPBA (Shattuck et al., 2008), and IXI (Avants et al., 2009), and one public abdomen CT dataset FLARE (Ma et al., 2022). We preprocess these four datasets as follows. For the Mindboggle dataset, we choose the NKI-RS-22 and NKI-TRT-20 sub-datasets from Mindboggle101 as the training set and OASIS-TRT-20 as the test set, following the prior study (Ma et al., 2024). We crop each volume to $160\times 192\times 160$ with an isotropic voxel size of 1 mm, while all images are pre-aligned with the MNI 152 template space. The LPBA dataset contains 40 T1-weighted MRI volumes, each annotated with 54 manually labeled Regions of Interest (ROIs). All volumes have been rigorously pre-aligned to the MNI 305. Specifically, each volume is cropped to $160\times 192\times 160$ voxels with an isotropic voxel size of 1 mm. We use 31 volumes for training and 9 for testing, following the prior study (Wang et al., 2024). The IXI dataset contains the original brain MRI images provided by (Fischl, 2012) and their corresponding 3D volume data after affine registration. All volumes are resampled to a voxel size of $80\times 96\times 112$ (isotropic voxel size of 2 mm), and for training, we use 407 3D volumes (total $407\times 406$ paired samples), and for testing, 58 volumes (total $58\times 57$ paired samples), following the prior study (Wang et al., 2024). The FLARE dataset consists of 361 scans annotated with four organ labels, namely liver, kidney, spleen, and pancreas. Each volume is cropped to $128\times 128\times 96$ voxels with an isotropic voxel size of 2.5 mm. We allocate 341 volumes for training and 20 for testing following the (Ma et al., 2024).
@@
Metrics: To comprehensively assess the performance of various registration models, we adopt multiple quantitative metrics. The Dice score (Dice, 1945) measures region‑wise overlap, while the 95% maximum Hausdorff Distance (HD95) (Huttenlocher et al., 1993) and the Average Symmetric Surface Distance (ASSD) (Taha and Hanbury, 2015) quantify the similarity of the regional profiles. The quality of the estimated deformation field $\varphi$ is assessed by the percentage of voxels with non-positive Jacobian determinant $|{J}_{\varphi}|_{\leq 0}$↓ (i.e., folded voxels). In addition, we report Mean Squared Error (MSE), Inference Time (s), and model size (the number of parameters) for a comprehensive performance analysis.
@@
Implementation Details: Our Model is implemented using Pytorch and trained on an NVIDIA GeForce RTX 4090 GPU. Following the prior study (Ma et al., 2024), we normalize all images to the range $[0,1]$, optimize using the ADAM optimizer (Kingma, 2014) at a learning rate of $1\times 10^{-4}$, and set the size of volumetric patch $n$, hyperparameter $\lambda$, and batch size to 27, 1, and 1, respectively. In FARM, we adopt a reduction ratio $r=16$. In TCI, we set stability threshold $\delta_{s}$, convergence threshold $\delta_{c}$, the window size $t$, and the maximum number of iterations $k_{max}$ to 0.005, 0.005, 3, and 10, respectively.
Appendix B Visualization of Feature Attention Patterns for AP and RP
Figure B: Visualization of Feature Attention Patterns for AP and RP.
As shown in Figure B, we illustrate the distinct feature attention patterns of AP and RP at each decoding Layer. The AP heatmaps concentrate more strongly on core registration regions (e.g., central gray matter), exhibiting a markedly higher response density, whereas the RP responses are more uniformly distributed. In addition, at deep decoder layers (e.g., $4$th Layer, corresponding to the coarse registration stage), large global structural discrepancies between the warped and fixed images encourage the network to exploit multi-regional features, resulting in widespread high-response regions in the heatmaps. As decoding proceeds from deep to shallow layers, registration errors decrease and anatomical structures become increasingly aligned, so the model primarily focuses on local regions for fine-grained adjustments. These visualizations demonstrate the rationality of our dual-pathway design.
Appendix C Performance Comparison on HD95, ASSD, and MSE Metrics
@@
Table C: Performance comparison of different registration models on the Mindboggle, LPBA, and IXI datasets evaluated by HD95, ASSD, and MSE metrics
@@
HD95↓ ASSD↓ MSE($10^{-3}$)↓ HD95↓ ASSD↓ MSE($10^{-3}$)↓ HD95↓ ASSD↓ MSE($10^{-3}$)↓
U-Net VM (CVPR’18) 5.683 1.726 7.449 6.557 2.089 2.293 1.460 0.584 0.727
@@
NAP (BSPC’25) 4.832 1.373 4.031 5.570 1.384 0.707 1.228 0.470 0.233
[…]
In Table C, we present the performance of different registration models on the Mindboggle, LPBA, and IXI datasets, evaluated by the HD95, ASSD, and MSE metrics. The experimental results demonstrate that iPEAR outperforms the other registration models in most scenarios and achieves SOTA results on at least two metrics for each dataset.
Appendix D Visualization of Registration Results on the LPBA Dataset
Figure D: Visualization of registration results and corresponding deformation fields across different models on the LPBA dataset.
In Figure D, we present the registration results and corresponding deformation fields across different models on the LPBA dataset. On LPBA, performance differences among models are more noticeable in the structural region of the brain images, highlighted using the blue box. These visualizations further show that, after warping with the deformation fields estimated by iPEAR, the moving images exhibit closer alignment with the fixed images.
Appendix E Quantitative Analysis on Different Anatomical Structures
Figure E.1: Performance comparison of different models in seven regions of the LPBA dataset.
Figure E.2: Performance comparison of different models in seven regions of the Mindboggle dataset.
Figure E.3: Performance comparison of different models in thirty regions of the IXI dataset.
Following the prior study (Wang et al., 2024), we evaluate iPEAR’s registration performance on selected regions in both hemispheres to assess its accuracy across diverse anatomical structures. Specifically, Figure E.1 presents the Dice scores for seven brain regions in the LPBA dataset, including the frontal, parietal, temporal, occipital, cingulate, hippocampus, and putamen. Figure E.2 presents the Dice scores for seven brain regions in the Mindboggle dataset, including the frontal, parietal, temporal, occipital, third ventricle, fourth ventricle, and brainstem. Figure E.3 presents the Dice score across 30 regions of the IXI dataset. All these experimental results demonstrate that iPEAR outperforms existing registration models.
Appendix F Performance Comparison of Different Iteration Strategies
Table F: Performance comparison of different iteration strategies on the Mindboggle and FLARE datasets
Mindboggle FLARE
Dice(%)↑ MSE($10^{-3}$)↓ Iter Dice(%)↑ MSE($10^{-3}$)↓ Iter
IIRP 65.8 4.16 [1.2, 3.5, 2.0, 2.0] 73.8 3.59 [1.3, 3.4, 3.0, 2.6]
IIRP (Iter = 2) 65.8 4.15 [2.0, 2.0, 2.0, 2.0] 73.6 3.66 [2.0, 2.0, 2.0, 2.0]
IIRP (Iter = 3) 66.0 4.03 [3.0, 3.0, 3.0, 3.0] 74.1 3.50 [3.0, 3.0, 3.0, 3.0]
IIRP (Iter = 4) 65.5 4.49 [4.0, 4.0, 4.0, 4.0] 73.7 3.61 [4.0, 4.0, 4.0, 4.0]
[…]
In this section, we examine the effectiveness of TCI in mitigating premature termination, and report the number of iterations determined by TCI and by the single-stage stopping strategy, respectively. As shown in Table F, on the Mindboggle and FLARE datasets, TCI judiciously increases the average number of iterations across all decoding levels, with the largest increase at $4$th Layer. The average number of iterations rises from 1.2 to 3.9 and from 1.3 to 3.0 on the two datasets, respectively. This deeper iterative refinement reduces the risk of early-stage under-optimization and helps limit the accumulation of anatomical misalignments. In addition, we compare TCI with a fixed-iteration setting, and the results show that TCI yields better registration performance than using a fixed number of iterations.
Appendix G Effectiveness of TCI at Each Decoding Layer
Table G: Ablation study on the effectiveness of TCI on various decoding layers
@@ Appendix G Effectiveness of TCI at Each Decoding Layer
66.0 73.1 79.9
66.2 73.4 80.0
67.3 73.6 80.2
67.9 73.7 80.6
In this section, we conduct ablation studies to assess the effectiveness of TCI at each decoding layers. Omitting TCI at a given layer fixes the number of iterations to one, thereby preventing iterative refinement of the deformation field. As shown in Table G, progressively enabling TCI across additional layers steadily improves iPEAR’s performance. Notably, applying TCI at higher-resolution layers (e.g., Layer 1) yields larger performance gains than at lower-resolution layers (e.g., Layer 4). A plausible reason for this is that higher-resolution feature maps more closely match the original image dimensions, preserving fine-grained structural details. Consequently, determining the number of iterations for higher-resolution layers yields greater performance improvements.
Appendix H Impact of Varying Values of Thresholds
Table H: Ablation study on thresholds $\delta_{s}$ and $\delta_{c}$ values
@@ Appendix H Impact of Varying Values of Thresholds
$0.01$ 65.9 66.3 73.1 73.2 80.4 80.4
$0.005$ 66.8 67.9 73.4 73.7 80.3 80.6
$0.001$ 67.0 67.6 73.3 73.4 80.3 80.2
Figure H.1: Statistics of the number of iterations across different decoding layers.
In this section, we investigate the sensitivity of the predetermined thresholds $\delta_{s}$ and $\delta_{c}$. As shown in Table H, these two thresholds exhibit low sensitivity to variation and have minimal impact on the overall performance of iPEAR. Nonetheless, we choose the middle range values, i.e., $\delta_{s}=0.005$ and $\delta_{c}=0.005$, which yield the best performance
In addition, in Figure H.1, we present the number of iterations per decoding layer determined by TCI on the IXI dataset. The experimental results demonstrate that three or four iterations suffice for convergence in most cases. Hence, our predefined maximum number of iterations $k_{max}=10$ is reasonable.
Appendix I Impact of Different TCI Similarity Metrics
In this section, we investigate how three similarity metrics, namely Normalized Cross-Correlation (NCC), Mean Absolute Error (MAE), and Mean Squared Error (MSE), affect the performance of TCI on the IXI dataset. As shown in Table I, NCC delivers the best results. Hence, we adopt NCC as the similarity metric for all experiments in this paper.
Table I: Ablation study on the effectiveness of different similarity metrics adopted in TCI
@@ Appendix I Impact of Different TCI Similarity Metrics
Appendix J Impact of Different Sliding Window Sizes
Table J: Ablation study on the different window sizes
IXI Mindboggle
$t$ Dice(%)↑ HD95↓ ASSD↓ MSE($10^{-3}$)↓ Time(s)↓ Dice(%)↑ HD95↓ ASSD↓ MSE($10^{-3}$)↓ Time(s)↓
3 80.6 1.23 0.46 0.23 0.12 67.9 4.91 1.37 3.97 0.19
4 80.9 1.22 0.46 0.21 0.15 67.4 4.93 1.38 3.89 0.22
5 80.7 1.27 0.49 0.23 0.29 67.0 4.94 1.39 4.06 0.35
In Table J, we present the effect of window size $t$ on the performance of iPEAR. On the IXI dataset, increasing $t$ from 3 to 4 yields marginal performance gains but incurs longer inference time. Increasing $t$ further to 5 degrades performance across all metrics, suggesting that $t=4$ provides sufficient iterative refinement without incurring excessive iterations. On the Mindboggle dataset, setting $t=3$ outperforms $t=4$ on most metrics, likely because $t=4$ induces excessive iterations in TCI’s deformation refinement. Therefore, we set $t=3$ for all experiments in this paper.